OpenStudy (anonymous):

Solve by graphing. x2 + 2x – 3 = 0

OpenStudy (anonymous):

Well if you have a calculator you could certainly graph it. But just factor it instead! No need for the quadratic formula here.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I assume you have to graph this by hand, so get out a sheet of paper (if you haven't already) and start plotting points. I suggest using only integers as x since your answer will be an integer. The answer should become clear as you graph one point at a time

OpenStudy (anonymous):

so what do you need help with? are you selecting one of those graphs?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you eliminate any of them as being impossible?

OpenStudy (anonymous):

Check your answer. Do the values of x beside the third graph make the equation true?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

Oh I see the problem. The third graph is right, but the x-values beside it are wrong.

OpenStudy (anonymous):

it would be -1,3 so is that right when I plugged that it for x I didn't get 0 but I might not be doing it right

OpenStudy (anonymous):

no it would be 1,-3

OpenStudy (anonymous):

so 1*1=1+ 2*-3=-6 - 3=0 but it doesn't it equals -9

OpenStudy (anonymous):

none of them are going to be equal to 0 as far as I can tell

OpenStudy (anonymous):

\[x ^{2}+2x-3=0\] test x=1,-1,3,-3\[1*1+2-3=3-3=0\]\[(-1*-1)-2-3=-1-3=-4\]\[3*3+6-3=12\]\[(-3*-3)-6-3=9-9=0\]

OpenStudy (anonymous):

I'll help you with the factor :) (x+3)* (x-1) = 0. This means that x is 0 at -3, and at 1. Which plots have x being 0 at -3 and at 1? That leaves us graphs A & C. We know the function is positive so that leaves C as the correct answer.

OpenStudy (anonymous):

You are so welcome.